OpenStudy (anonymous):

x2 − 2xy + y2 + 2x + 2y = 0. factor

OpenStudy (anonymous):

x'2 = 8y' – 2 y'2 = 4x x2 = 4y − 2 y'2 = 2x

OpenStudy (anonymous):

What would the equation of the parabola be in terms of the rotated axes?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

I chopped the question up so people would come

OpenStudy (anonymous):

@EclipsedStar

OpenStudy (anonymous):

n engineer designing a roller coaster is using an advanced software application to test the strength of the structure. A part of the structure has a parabolic shape with the equation x2 − 2xy + y2 + 2x + 2y = 0. However, the software doesn't recognize any equation involving an xy term. What would the equation of the parabola be in terms of the rotated axes?

OpenStudy (michele_laino):

please wait I try

OpenStudy (anonymous):

yea take your time

OpenStudy (michele_laino):

here, we have the subsequent matrix equation: \[\left( {\begin{array}{*{20}{c}} x \\ y \end{array}} \right) = \left( {\begin{array}{*{20}{c}} {1/\sqrt 2 }&{1/\sqrt 2 } \\ { - 1/\sqrt 2 }&{1/\sqrt 2 } \end{array}} \right)\left( {\begin{array}{*{20}{c}} {x'} \\ {y'} \end{array}} \right)\] where x' and y' are the new coordinates. Namely we have to compute the eigenvalues and the eigenvectors of the quadratic part of your parabola

OpenStudy (michele_laino):

so we can write: \[\begin{gathered} x = \frac{{x' + y'}}{{\sqrt 2 }} \hfill \\ y = \frac{{x' - y'}}{{\sqrt 2 }} \hfill \\ \end{gathered} \] those are the requested equation of our rotation

OpenStudy (michele_laino):

Now we have to substitute those equations into your parabola, in order to rewrite your parabola in term of x' and y',so what do you get?

OpenStudy (anonymous):

substitute \[x=\frac{ x'+y }{ \sqrt{2} } and y=\frac{ x' -y'}{ \sqrt{2} }\] x2 − 2xy + y2 + 2x + 2y = 0

OpenStudy (anonymous):

that?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

x = 1/4, y = -3/4

OpenStudy (anonymous):

you wont me to substitute the factor in to the equation

OpenStudy (anonymous):

the solution?

OpenStudy (michele_laino):

for example: \[\large \begin{gathered} {x^2} + {y^2} - 2xy + 2x + 2y = \hfill \\ \hfill \\ = {\left( {\frac{{x' + y'}}{{\sqrt 2 }}} \right)^2} + {\left( {\frac{{x' - y'}}{{\sqrt 2 }}} \right)^2} - 2\left( {\frac{{x' + y'}}{{\sqrt 2 }}} \right) \cdot \left( {\frac{{x' - y'}}{{\sqrt 2 }}} \right) + \hfill \\ + 2\frac{{x' + y'}}{{\sqrt 2 }} + 2\frac{{x' - y'}}{{\sqrt 2 }} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

I got this: \[\Large y{'^2} + \sqrt 2 x' = 0\] please check my computations

OpenStudy (anonymous):

y'^2 = 2x?

OpenStudy (michele_laino):

I think: y'2=-sqrt(2)x'

OpenStudy (anonymous):

so its D

OpenStudy (michele_laino):

please can you check your options?

OpenStudy (anonymous):

It is D:)

OpenStudy (michele_laino):

ok!